OpenStudy (anonymous):

You drive a car that holds 15 gallons of gasoline. Your car uses 0.03 gal per mi. Which equation best represents this information?

OpenStudy (anonymous):

"which" implies some kind of choices...

OpenStudy (anonymous):

A. y = 0.03x + 15 B. y = 0.03x - 15 C. y = -0.03x + 15 D. y = -0.03x - 15

OpenStudy (radar):

It deserves to be closed.

OpenStudy (radar):

What is y to represent? What is x to represent etc.

OpenStudy (anonymous):

Think about this: you start off with 15 gallons, and as time passes, you have LESS by 0.03. Therefore, the y-int would be (0,15) because that is how much you have at time 0 (start), and you have a slope of -0.03 because you lose that much per mile. Now, can you tell me which equation that would be?

OpenStudy (radar):

In spite of it all, I think B is the best ansnwer provided y represents the amount of gas remaining in tank after x miles.

OpenStudy (radar):

Is that what you're thinking the cholcluver445?

OpenStudy (anonymous):

Wouldn't it be C? You have a neg slope because you lose gas as you drive more.

OpenStudy (anonymous):

 Think of it like this.